Integrated Nanosystems Research Facility

This weeks updates:
  • STS ICP is down due to a communication error. We’ve contacted STS and are waiting to hear back from them. The problem seems to be a bad module that will not communicate with the main PC.
  • Replaced the UPS system on the Cannon Stepper. System is now working within normal perimeters.
  • Finished repairing the CHA Evaporator Cryo Compressor and regenerated the system. Will do a dummy in the next day or so.
  • Allocated time to the FEI E-beam writing system. Reconfigured the computer cables so system is being checked out to see if the Nabady software will communicate with the stage.
  • The Gasonics asher is back online. The end of point error was corrected and is working within normal perimeters.
  • BMR Nitrite Oxide system, the gas manifold needs cleaned due to deposition being ununiformed and deposition rate has dropped.